hi, i would like to solve this problem step by step to find X1.
9X1 + 6(200-X1) = 1566
i know the answer that X1 = 122 but i don't know how to go over the step inorder to get that answer, can you help me =)

Answer by funmath(2933)

solve this problem step by step to find X1.
9X1 + 6(200-X1) = 1566
9x1+6(200)+6(-x1)=1566
9x1+1200-6x1=1566
9x1-6x1+1200=1566
(9-6)x1+1200=1566
3x1+1200=1566
3x1+1200-1200=1566-1200
3x1+0=366
3x1=366
3x1/3=366/3
x1=122
:
Happy Calculating!!!